OpenStudy (kitkat16):

Help please! If the vertex of a parabola is (-1.5, -12.5), and its y-intercept is (0, -8). What are the x-intercepts of the parabola? anyone?

OpenStudy (kitkat16):

@lsherron can you help please?

OpenStudy (anonymous):

0 is the x int and -8 is the y int

OpenStudy (kitkat16):

(0,-8) is the y intercept

OpenStudy (kitkat16):

The x intercepts are 2 of these below (-2,0) (-1,0) (-4,0) (2,0) (2,0) (1,0) (4,0) (-2,0)

OpenStudy (anonymous):

(-2,0)

OpenStudy (kitkat16):

for both of them? I have to pick two

OpenStudy (kitkat16):

can I write this into an equation to check the vertex?

OpenStudy (anonymous):

yes

OpenStudy (kitkat16):

can u show me

OpenStudy (anonymous):

ok

OpenStudy (kitkat16):

thanks

OpenStudy (kitkat16):

@TheSmartOne

OpenStudy (kitkat16):

@jim_thompson5910 can you help me please?

OpenStudy (anonymous):

If we write our equation in completed square form using coefficients a and b, which we want to find, then we have this:\[y=(x+a)^2+b\]Using the two coordinates we have: (-1.5, -12.5) and (0, -8):\[-8=(0+a)^2+b\]\[-8=a^2+b\] \[-12.5=(-1.5+a)^2+b\]\[-12.5=2.25-3a+a^2+b\]Substituting what we found above into this:\[-12.5=2.25-3a+(a^2+b)\]\[-12.5=2.25-3a-8\]\[3a=6.75\]\[a=2.25\] Using this in the equation: \[-8=a^2+b=2.25^2+b=5.0625+b\]\[b=-13.0625\] Giving us the answer: \[y=(x+2.25)^2-13.0625\]

OpenStudy (kitkat16):

thank you but does this show me the x intercepts?

OpenStudy (anonymous):

Oops sorry, misread the question. The x intercepts are when y=0, so set y=0 and solve for x. I'll help if you get stuck, but I imagine you can do this.